# Analyzing Logs of Large-Scale Software Systems using Time Curves Visualization

### v1

@@ Abstract
Logs are crucial for analyzing large-scale software systems, offering insights into system health, performance, security threats, potential bugs, etc. However, their chaotic nature—characterized by sheer volume, lack of standards, and variability—makes manual analysis complex. The use of clustering algorithms can assist by grouping logs into a smaller set of templates, but lose the temporal and relational context in doing so. On the contrary, Large Language Models (LLMs) can provide meaningful explanations but struggle with processing large collections efficiently. Moreover, representation techniques for both approaches are typically limited to either plain text or traditional charting, especially when dealing with large-scale systems. In this paper, we combine clustering and LLM summarization with event detection and Multidimensional Scaling through the use of Time Curves to produce a holistic pipeline that enables efficient and automatic summarization of vast collections of software system logs. The core of our approach is the proposal of a semimetric distance that effectively measures similarity between events, thus enabling a meaningful representation. We show that our method can explain the main events of logs collected from different applications without prior knowledge. We also show how the approach can be used to detect general trends as well as outliers in parallel and distributed systems by overlapping multiple projections. As a result, we expect a significant reduction of the time required to analyze and resolve system-wide issues, identify performance bottlenecks and security risks, debug applications, etc.
@@ I Introduction
Figure 1: Example Time Curve corresponding to the analysis of the execution of a stream processing framework. Note that the dashed-gray line is not produced by the method itself but was added to highlight a characteristic of the projection.
Several sophisticated and automated solutions exist for tasks such as anomaly detection [4] or failure prediction [5], which have proven that log data can be useful. However, exploring and understanding the system log still remains time consuming and not trivial. Therefore, in this paper we tackle a more general problem – the lack of explainability in log data analysis. In particular, we explore the potential of the visualization techniques to explain the evolution of a computing system, purely based on the log data, and suggest a new approach to solving the problem. This is showcased in Fig. 1, which corresponds to the automatic analysis of the execution of a stream processing framework [6].
Each node (represented with vertices) corresponds to an event state in the application, and the proximity between nodes depicts how similar they are, i.e., closer nodes include more similar sets of logs. While specific details about the analysis are given in Section IV-A, numerous visual patterns may be observed from the visualization itself, such as (1) the relative symmetry along the dashed-gray line, (2) the cyclic nature of the two halves, (3) the clusters of repetitive states (shown as halos), etc.
We propose a holistic pipeline that enables to automatically extract and visualize the main patterns of the evolution of a system and to identify events of interest based exclusively on the underlying logs. Therefore, our approach can result in a significant reduction of the time required to analyze system outages, identify performance bottlenecks, perform root cause analysis, etc., thus benefiting a wide spectrum of the industry.
@@
    Several extensions to the Time Curves [7] technique, including the comparison of multiple systems at once, animating curves or enriching nodes via LLM summarization.
@@
    The experimental approach as a validation and reference method for the analysis of log datasets. This is accomplished by using a controlled experiment, a public dataset and a performance benchmark.
@@
    The experimental analysis from which we gather insights about the log datatsets and can be reused and contrasted with further analyzes.
The remainder of this paper is structured as follows: Section II discusses the current state-of-the-art of log analysis and visualization; Section III introduces the concept of Time Curves and the components of the processing pipeline, such as event detection, clustering and distance computation; Section IV describes the experiments performed to depict the capabilities of the method. Lastly, Section V concludes this study and discusses future work.
@@ II Related work
The most common tools for visualizing log data are dashboards, which typically display numerous detailed histograms and line plots simultaneously, allowing for interactive navigation (e.g., selecting timeframes or applying filtering queries). While dashboards are theoretically well-suited for aiding in problem analysis, mastering them can be challenging, and advanced interactions require significant time investment. Consequently, dashboards are generally not intended for use by a non-expert audience. However, an overall analysis of system evolution can be valuable even for individuals with limited knowledge of the system, such as small online service owners, salespeople, client support staff, and developers from related domains. This underscores the need for simple and intuitive visualization strategies.
@@
For instance, Streamgraphs [8] or stacked bar charts can show composition over time (counts of similar logs over time), but require categorization of the data into distinct disjoint classes with certain properties of interest, which in the case of logs is not trivial since the most valuable information in the system log typically lies in the semantic meaning of log messages, which is hardly measurable.
@@
Such lack of intuitive features can potentially be addressed through the use of embeddings. Visualizing the embedding space can provide insights into the composition of semantic content by training or utilizing pre-trained text embeddings and visualizing them, such as through down-projections on scatter plots [9]. While a robust embedding transformation can effectively display the distribution of log messages and their semantic similarities, it struggles to illustrate the evolution of the system state due to the loss of temporal context. To incorporate chronological order, one can adapt trajectory-based methods similar to those in [10]. However, in their approach, encodings need to be manually designed, which can be a challenging task depending on the data type.
@@
Even with effective trajectory bundling as suggested in [10], the results can become visually complex. In practice, system logs can easily reach millions of lines and gigabytes of data. Therefore, grouping consecutive log messages into events is crucial not only for visualization purposes but also due to computational limitations. For example, feeding an entire log file directly into any large language model (LLM) can be either unfeasible or overly expensive [3]. While utilizing the context window may suffice for tasks such as anomaly detection [11], it may be unsuitable for providing a comprehensive overview of system evolution. To enable AI assistance in analyzing the evolution of the system state, we need to develop a more abstract and concise representation of the log messages that still preserves all essential semantic information contained in the original raw data.
@@
Thinking in terms of groups of log records (referred to as events) allows for the consideration of visually simpler techniques, such as timelines. Timelines are easy to interpret and effectively display the flow of events over time. However, they do not incorporate semantic similarity, making state analysis challenging. One might attempt to extend timelines to 2-D graphs or utilize various bump charts [12]. This approach, however, requires splitting the sequence into meaningful sub-parts based on content, as demonstrated with categories of user actions [13]. Unfortunately, it is not directly possible for the case of log records, since in order to divide the monitored processes into relevant groups and assign each log record to one of such groups, extensive knowledge about the events of the system is required. But this knowledge is precisely what the proposed method attempts to extract and present.
@@
Our proposal is to integrate a timeline of events with a scatter plot that projects the corresponding semantic content based on similarity, computed directly without the need for embeddings or feature design. The goal is to visualize both the temporal and similarity dimensions of the system’s various states within a single plot. A direct combination of these techniques is the Time-connected scatter plot [14], initially used in journalism. A more effective method for encoding time than static labels near data points is the use of colorful segments with discretized time, as introduced by Time-segmented scatter plots [15]. However, connecting points with straight lines can result in high visual complexity. Therefore, in this manuscript, we consider Time Curves [7] as a more promising approach. Time Curves produce smooth curves where the connection between data points are easier to track and interpret while still allowing for the analysis of the evolution of the system state.
@@ III Visualizing Logs with Time Curves
The proposed method is designed to aggregate information from a large collection of system logs into a single and interactive representation. Such representation uses the Time Curves technique to show a series of important events in a 2D space based on the given distance. The data points in the plot correspond to system states or events, which will be spatially closer if their content (i.e., the log messages produced during the event) is also similar. This enables both to analyze single log collections and to overlap concurrent system executions to show overall trends and outliers. Furthermore, the representation can be enriched with domain-specific information or summarizations provided by LLMs.
@@ III-A Introduction
In the original Time Curves paper [7], the authors described several use cases, including analyzes of document changes (such as in Wikipedia article revisions) or video footage. In this paper, the data consists of system logs, where each item is a plain text message, typically including a timestamp along with optional tags and a message payload. While it might appear to be similar to the article revision use case, the sheer cardinality of logs makes it hardly comparable. System logs are often counted in the hundreds of thousands or millions, which makes a projection of such a high number of data points impractical to say the least. Therefore, our proposal must cope with an increase in dimensionality: instead of directly projecting logs as points for Time Curves, it becomes necessary to first aggregate logs into sets (or events).
@@ III-B Workflow overview
The method is composed of (1) data preprocessing, (2) grouping and template extraction, (3) distance computation and (4) projection and summarization. The first stage focuses solely on parsing the input logs (timestamp and other optional tags), joining multi-line logs together and forming log records. The second stage breaks down the preprocessed log records into a sequence of events by grouping records in discrete checkpoints (points in time associated with part of the data) and mapping records to templates by performing clustering. The third stage consists of computing distances between templates and aggregating them to get distances between checkpoints (represented as sets of templates). The last stage completes the method by projecting checkpoints onto a 2-dimensional plane according to computed distances and joins them in chronological order to produce a Time Curve) while also enabling annotations and enrichment via LLMs.
@@ III-C Preprocessing
As explained in the background, the main difference between a Time Curve and a similarity-scatter projection is the fact that the time dimension is kept. Therefore, in the case of log data, we need to extract time information. In addition, we also parse other useful information such as the severity level or particular keywords (e.g., Exception).
While there is no single logging standard, most common logging formats usually include a timestamp as well as a tag representing the severity level before including the payload, e.g.: ‘‘23-09-12 13:01 WARN - Unexpected value ...’’.
@@
The severity level is extracted by matching keywords such as WARN or INFO and mapping them to integer constants. The higher the mapping values, the higher the impact they have on the grouping decisions into events. In order to treat the data as a proper time series of log records, we sort records in ascending order according to the timestamp and therefore do not use the original ordering which might be tampered by batching or bulking mechanisms.
@@
Once records have been properly parsed, the next step is to group them into events, i.e., split the complete collection of logs into multiple checkpoints according to detected events.
III-D Event grouping and clustering
@@
Event detection (or grouping into checkpoints) is an inherently difficult problem. Multiple time series analysis or change detection algorithms exist in the literature (e.g., see [16] and [17]). Nonetheless, they are typically not designed for discrete $n$-ary valued time series detection, such as in the case of analyzing severity level. For instance, when an error occurs, multiple other errors might follow, or a series of different types of records might be printed. The grouping approach must be able to consider such scenario and group all corresponding logs into a single event.
@@
While prior information about the logging system may be used it to improve the final grouping, the default method takes into account the severity level of log record, the time gaps in between records and the size limits of checkpoints.
@@
First, it sets the maximum number of log records in a checkpoint, which is estimated based on the desired number of checkpoints (generally between few tenths to a hundred). Then, it computes the time gaps between consecutive records and determines the positions of the $k$-largest gaps, where $k$ is configurable. If the $(k+1)$-th largest gap is equal to the $k$-th, $k$ is decremented until the gaps are no longer equal or $k$ reaches 0, indicating that more than $k$ largest gaps are equal and time gaps will be ignored.
@@
Next, the method identifies points of severity level change. To ensure robustness, it smooths the extracted severity levels by applying a convolution with a normalized asymmetric window. The decaying and smoothing windows cause the function value to increase rapidly with errors but require several error-free log records to return to the constant level. Such level should vary depending on the system itself but is set by default to the mapping value of INFO plus a small constant. Event change points are generated both when the function raises above the threshold (indicating the start of an event) and when it returns below the threshold (indicating the end of an event).
@@
All this information is then combined. The process starts by creating a checkpoint with only the first log record. Before adding each next log record, the method checks whether the maximum checkpoint size is reached, whether a severity level change point occurs or if the current time gap is within the list of the $k$-largest time gaps. If any of these conditions are met, a new event checkpoint is created, and log records are added accordingly. This process repeats until the last record is added to a checkpoint. The outcome is as follows:
@@
    Errors always trigger an event unless they are part of a succession of previous errors occurred in the near past, in which case they get bundled together.
@@
    An excessive number of records produced within a small time frame triggers an event.
After records have been grouped into checkpoints, a clustering algorithm [18] is applied to extract templates. Templates are composed of static parts (such as service names) and variable parts (such as IP addresses). Log records are then mapped to their corresponding templates and the variable parts are masked, thus reducing the cardinality at the expense of exactitude. This is due to (1) sequence order is lost within checkpoints and (2) the masked variable parts may have contained important information. However, without this step, the sheer number of log records would make the projection useless. In general, this results into a coarse-grain aggregation of the major events and changes that happened within the system. In case a finer-grain analysis is required (e.g., small datasets), the size of the checkpoints can be reduced to as small as only one record, thus reducing the amount of information lost due to clustering. Once clustering has been applied, each checkpoint can be represented by a set of all the unique log templates which it contains.
@@ III-E Distance properties
    Checkpoints containing similar templates should have a small distance. Therefore, the distance should account for variations between templates. For instance, using the Jaccard distance [20] would result in large distance between two sets containing System started rotating logs and System finished rotating logs, despite the fact that their semantics are similar since they describe the state of the rotation of logs.
@@
III-F Notation
We use $P$ to denote the Time Curve as an ordered set of all $N$ checkpoints, meaning $P=\big{(}P_{0},P_{1},...,P_{N-1}\big{)}$. Checkpoints are denoted either as $i$-th element of $P$ such as $P_{i}$ or explicitly with lower-case letters (e.g., $x,y,z$). Note that checkpoints are also sets of the representative log templates, e.g., checkpoint $x$ (with $n$ templates) is denoted as $x:=\{x_{0},\dots,x_{n-1}\}$, where all $x_{i}$ are string log templates.
||$\cdot$|| ⋅ | denotes cardinality of sets and for strings - the length in characters, e.g., $|x|:=n$ – number of templates in checkpoint $x$, $|x_{i}|$ – number of characters in string template.
III-G Distance computation
For similarity between elements we chose the Levenshtein distance, because it is standard for string comparison and there are multiple computational optimizations available. However, any distance function designed for variable-length strings can be used as well, e.g., the Longest Common Subsequence or the $q$-gram [21] distances.

### v2

@@ Abstract
Logs are crucial for analyzing large-scale software systems, offering insights into system health, performance, security threats, potential bugs, etc. However, their chaotic nature—characterized by sheer volume, lack of standards, and variability—makes manual analysis complex. The use of clustering algorithms can assist by grouping logs into a smaller set of templates, but lose the temporal and relational context in doing so. On the contrary, Large Language Models (LLMs) can provide meaningful explanations but struggle with processing large collections efficiently. Moreover, representation techniques for both approaches are typically limited to either plain text or traditional charting, especially when dealing with large-scale systems. In this paper, we combine clustering and LLM summarization with event detection and Multidimensional Scaling through the use of Time Curves to produce a holistic pipeline that enables efficient and automatic summarization of vast collections of software system logs. The core of our approach is the proposal of a semimetric distance that effectively measures similarity between events, thus enabling a meaningful representation. We show that our method, based on logs collected from different applications, can explain the behavior of a system over time without prior knowledge. We also show how the approach can be used to detect general trends as well as outliers in parallel and distributed systems by overlapping multiple projections. As a result, we expect a significant reduction in the time required to analyze system-wide issues, identify performance bottlenecks and security risks, debug applications, etc.
@@ I Introduction
Figure 1: Example Time Curve corresponding to the analysis of the execution of a stream processing application, where data points represent events of the system, consecutive events in time are connected, while similar events are positioned close. Color of connection shows time, i.e. start of observation purple – end of observation green.
Several sophisticated and automated solutions exist for tasks such as anomaly detection [4] or failure prediction [5], which have proven that log data can be useful. However, in practice, users typically examine system logs only after an automated alert is triggered, either because the issue cannot be resolved automatically or because they seek a deeper understanding of the problem. In both scenarios, the lack of explainability complicates log analysis. At Dynatrace, the immense volume of log ingestion and the need for real-time analytics have motivated us to explore visualization techniques that can summarize and explain the evolution of a software system using only log data. We believe that the suggested visualization can serve as an effective entry point for analyzing software systems by providing a comprehensive overview of system behavior and facilitating quick insights, thus enabling efficient exploration. Initially, one can quickly identify visual patterns and easily review underlying logs. If further investigation is needed, one can then dive into standard dashboards, integrating the insights gained (e.g., relevant time frame) with other available information (e.g., metrics).
Fig. 1 illustrates the Time Curve of the automatic analysis of the execution of a stream processing application [6] using the proposed method.
Figure 2: Workflow illustration of the proposed method to generate a Time Curve from a set of logs.
Each data point corresponds to a system event, and the proximity between points depicts similarity between the corresponding sets of logs of the events. Although specific details on the analysis are given in Section IV-A, some visual patterns from [7] can already be intuitively perceived from the visualization itself, such as clusters of events (shown as halos indicating presence of almost identical events) and alternations (representing the repetitive nature of the system behavior). Note that a glossary containing definitions of visual patterns and other terms used in the paper is available in [8].
We propose a holistic pipeline (see Fig. 2) that enables to automatically extract processing trends and visualize the evolution of a system based exclusively on the underlying logs. Therefore, our approach can result in a significant reduction of the time required to analyze system outages, identify performance bottlenecks, perform root cause analysis, etc., thus benefiting a wide spectrum of the industry.
@@
    Several extensions to the Time Curves [7] technique, including the comparison of multiple systems at once, animating curves or enriching them via LLM summarization.
@@
    The experimental approach as a validation and reference method for the analysis of log datasets. This is achieved by (1) using an experiment under controlled conditions, (2) a public dataset and (3) a performance benchmark.
@@
    The experimental analysis and insights gained from the log datasets, which can be compared with further experiments and results from other tools.
The remainder of this paper is structured as follows: Section II discusses the current state-of-the-art of log analysis and visualization; Section III introduces the concept of Time Curves and the components of the processing pipeline (Fig. 2), such as event detection, clustering and distance computation; Section IV describes the experiments performed to depict the capabilities of the method. Lastly, Section V concludes this study and discusses future work.
@@ II Related work
Dashboards are the most common tools for visualizing log data, typically featuring numerous detailed histograms and line plots that allow for interactive navigation, such as selecting time frames or applying filters. Although dashboards are theoretically well-suited for problem analysis, mastering them can be challenging, and advanced interactions require significant time investment. The extensive functionality of dashboards can overwhelm users and even cause confusion, especially among non-experts. However, an overall analysis of the evolution of the system can also be valuable for individuals with limited knowledge of the system, such as owners of small online services, salespeople, client support staff and developers from related domains. These points highlight the need for simple and intuitive visualization strategies.
@@
For instance, Streamgraphs [9] or stacked bar charts can show composition over time (counts of similar logs over time), but require categorization of the data into distinct disjoint classes with certain properties of interest. In the case of logs this is not trivial since the most valuable information typically lies in the semantic meaning of log messages, which is hardly measurable.
@@
Such lack of intuitive features can potentially be addressed through the use of embeddings. Visualizing the embedding space can provide insights into the composition of semantic content by training or utilizing pre-trained text embeddings and visualizing them, such as through down-projections on scatter plots [10]. While a robust embedding transformation can effectively display the distribution of log messages and their semantic similarities, it struggles to illustrate the evolution of the system due to the loss of temporal context. To incorporate chronological order, one can adapt trajectory-based methods similar to those in [11]. However, in their approach, encodings need to be manually designed, which can be a challenging task depending on the data type.
@@
Even with effective trajectory bundling as suggested in [11], the results can become visually complex. In practice, system logs can easily reach millions of lines and gigabytes of data. Therefore, grouping consecutive logs into events is crucial not only for visualization purposes but also due to computational limitations. For example, feeding an entire log file directly into any large language model (LLM) can be either unfeasible or overly expensive [3]. While utilizing the context window may suffice for tasks such as anomaly detection [12], it may be unsuitable for providing a comprehensive overview of the system over time. To enable AI assistance in analyzing the evolution of the system, we need to develop a more abstract and concise representation of the logs that still preserves all essential semantic information contained of the raw data.
@@
Thinking in terms of events (groups of consecutive logs) allows for the consideration of visually simpler techniques, such as timelines. Timelines are easy to interpret and effectively display the flow of events over time. However, they do not incorporate semantic similarity, making the analysis challenging. One might attempt to extend timelines to 2-D graphs or utilize various bump charts [13]. This approach, however, requires splitting the sequence into meaningful sub-parts based on content, as demonstrated with categories of user actions [14]. Unfortunately, it is not directly possible for the case of log records, since, in order to divide the monitored processes into relevant groups and assign each log record to one of such groups, extensive knowledge about the behavior of the system is required. However, this knowledge is precisely what the proposed method attempts to extract and present.
@@
Our proposal is to integrate a timeline of events with a scatter plot that projects the corresponding semantic content based on similarity, computed directly without the need for embeddings or feature design. The goal is to visualize both temporal and similarity dimensions of the system evolution within a single plot. A direct combination of these techniques is the Time-connected scatter plot [15], initially used in journalism. A more effective method for encoding time than static labels near data points is the use of colorful segments with discretized time, as introduced by Time-segmented scatter plots [16]. However, connecting points with straight lines can result in high visual complexity. Therefore, in this manuscript, we consider Time Curves [7] as a more promising approach. Time Curves produce smooth curves where the connection between data points are easier to track and interpret while still allowing for the analysis of the evolution of the system.
@@ III Visualizing Logs with Time Curves
The proposed method is designed to aggregate information from a large collection of system logs into a single and interactive representation. Such representation uses the Time Curves technique to show a series of data points in a 2D space based on the given distance measure. These data points in the plot correspond to system events, which will be spatially closer if their content (i.e., the logs) is similar. This enables both to analyze single log collections and to overlap concurrent system executions to show overall trends and outliers. Furthermore, the representation can be enriched with domain-specific information or summarizations provided by LLMs.
@@ III-A Introduction
In the original Time Curves paper [7], the authors described several use cases, including analyzes of document changes (such as in Wikipedia article revisions) or video footage. In this paper, the data consists of system logs, where each item is a plain text message, typically including a timestamp along with optional tags and a message payload. While it might appear to be similar to the article revision use case, the sheer cardinality of logs makes it hardly comparable. System logs are often counted in the hundreds of thousands or millions, which makes a projection of such a high number of data points impractical to say the least. Therefore, our proposal must cope with an increase in dimensionality: instead of directly projecting logs as data points for Time Curves, it becomes necessary to first aggregate the logs.
@@ III-B Workflow overview
The method (as shown in Fig. 2) is composed of (1) data preprocessing, (2) grouping and template extraction, (3) distance computation and (4) projection and summarization. The first stage focuses solely on parsing the input logs (timestamp and other optional tags), joining multi-line logs together and forming log records. The second stage breaks down the preprocessed log records into a sequence of events by grouping consecutive logs and mapping records to templates by performing clustering. The third stage consists of computing distances between templates and aggregating them to get distances between checkpoints (which represent events as sets of templates). The last stage completes the method by projecting checkpoints onto a 2-dimensional plane according to computed distances and joins them in chronological order to produce a Time Curve while also enabling annotations and enrichment via LLMs.
@@ III-C Preprocessing
The main difference between a Time Curve and a similarity-scatter projection is the fact that the temporal context is kept. Therefore, we extract timestamps from log messages. In addition, we parse other useful information such as severity level tags or particular keywords (e.g., Exception).
While there is no single logging standard, most common logging formats usually include a timestamp as well as a tag representing the severity level before including the payload, e.g., "23-09-12 13:01 WARN - Unexpected value ...".
@@
The severity level is extracted by matching keywords such as WARN or INFO and mapping them to integer constants. The higher the mapping values, the higher the impact they have on the grouping into events. In order to treat the data as a proper time series of log records, we sort records in ascending order of timestamps and therefore do not use the original ordering which might be tampered by batching or bulking mechanisms.
@@
Once records have been properly parsed, the next step is to group them into events, i.e., split the complete collection of logs into multiple chronologically consecutive sets.
III-D Event detection and clustering
@@
Event detection is an inherently difficult problem. Multiple time series analysis or change detection algorithms exist in the literature (e.g., see [17] and [18]). Nonetheless, they are typically not designed for discrete $n$-ary valued time series detection, such as in the case of analyzing severity level. For instance, when an error occurs, multiple other errors might follow, or a series of different types of records might be printed. The grouping approach must be able to consider such scenario and group all corresponding logs into a single event.
@@
While prior information about the logging system may be used it to improve the final grouping, the default method takes into account the severity level of log record, the time gaps in between records and the size limits of events.
@@
First, it sets the maximum number of log records for an event, which is estimated based on the desired number of data points (generally between a few tenths and a hundred). Then, it computes the time gaps between consecutive records and determines the positions of the $k$-largest gaps, where $k$ is configurable. If the $(k+1)$-th largest gap is equal to the $k$-th, $k$ is decremented until the gaps are no longer equal or $k$ reaches 0, indicating that more than $k$ largest gaps are equal and time gaps will be ignored.
@@
Next, the method identifies points of severity level change. To ensure robustness, it smooths the extracted severity levels by applying a convolution with a normalized asymmetric window. The decaying and smoothing windows cause the function value to increase rapidly with errors but require several error-free log records to return to the constant level. Such level should vary depending on the system itself but is set by default to the mapping value of INFO plus a small constant. Change points are detected both when the function raises above the threshold (indicating the start of an event) and when it returns below the threshold (indicating the end of an event).
@@
All this information is then combined. The process starts by creating an event with only the first log record. Before adding each next log record, the method checks whether the maximum event size is reached, whether a severity level change point occurs or if the current time gap is within the list of the $k$-largest time gaps. If any of these conditions are met, a new event is created, and log records are added accordingly. This process repeats until the last record is added to an event. The outcome is as follows:
@@
    Errors always trigger a new event unless they are part of a succession of previous errors occurred in the near past, in which case they get bundled together.
@@
    An excessive number of records produced within a small time frame triggers a new event.
After records have been grouped into events, a clustering algorithm [19] is applied to extract templates. Templates are composed of static parts (such as service names) and variable parts (such as IP addresses). Log records are then mapped to their corresponding templates and the variable parts are masked, thus reducing the cardinality at the expense of exactitude. This is due to (1) sequence order is lost within checkpoints and (2) the masked variable parts may have contained important information. However, without this step, the sheer number of log records would make the projection useless. In general, this results into a coarse-grain aggregation of the major changes that happened within the system. In case a finer-grain analysis is required (e.g., small datasets), the size of the events can be reduced to as small as only one record, thus reducing the amount of information lost due to clustering. Once clustering has been applied, each event can be represented as a checkpoint, i.e., by a timestamp and a set of all the unique log templates which it contains.
@@ III-E Distance properties
    Checkpoints containing similar templates should have a small distance. Therefore, the distance should account for variations between templates. For instance, using the Jaccard distance [21] would result in large distance between two sets containing "System started rotating logs" and "System finished rotating logs", despite their semantic similarities, i.e., both describe the rotation of logs.
@@
III-F Distance computation
We denote checkpoints with lower-case letters and templates with sub-indices, e.g., a checkpoint with $n$ templates is denoted as $x:=\{x_{0},\dots,x_{n-1}\}$, where all $x_{i}$ are string log templates. In addition, ||$\cdot$|| ⋅ | denotes the cardinality of sets and the length of strings, e.g., $|x|:=n$ is the number of templates in checkpoint $x$, $|x_{i}|$ is the number of characters in the string template.
The similarity between templates is computed via the Levenshtein distance, given its flexibility and available computational optimizations. However, any distance function designed for variable-length strings can be used as well, e.g., the Longest Common Subsequence or the $q$-gram [22] distances.